Expand Your Critical Thinking 25-02 a-c

Hawke Skateboards is considering building a new plant. Bob Skerritt, the company’s marketing manager, is an enthusiastic supporter of the new plant. Lucy Liu, the company’s chief financial officer, is not so sure that the plant is a good idea. Currently, the company purchases its skateboards from foreign manufacturers. The following figures were estimated regarding the construction of a new plant.


Cost of plant $3,040,000 Estimated useful life 15 years

Annual cash inflows 3,040,000 Salvage value $1,520,000

Annual cash outflows 2,690,000 Discount rate 11%


Bob Skerritt believes that these figures understate the true potential value of the plant. He suggests that by manufacturing its own skateboards the company will benefit from a “buy American” patriotism that he believes is common among skateboarders. He also notes that the firm has had numerous quality problems with the skateboards manufactured by its suppliers. He suggests that the inconsistent quality has resulted in lost sales, increased warranty claims, and some costly lawsuits. Overall, he believes sales will be $152,000 higher than projected above, and that the savings from lower warranty costs and legal costs will be $46,000 per year. He also believes that the project is not as risky as assumed above, and that a 9% discount rate is more reasonable.


Click here to view the factor table.


Answer each of the following.




Partially correct answer. Your answer is partially correct.  Try again.

   

Compute the net present value of the project based on the original projections. (If the net present value is negative, use either a negative sign preceding the number e.g. -45 or parentheses e.g. (45). Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 5,275.)


Net present value $Entry field with incorrect answer

(11,453,318))


Should the project be accepted?


The project should be Entry field with correct answer

rejected

.


LINK TO TEXT




Partially correct answer. Your answer is partially correct.  Try again.

   

Compute the net present value incorporating Bob’s estimates of the value of the intangible benefits, but still using the 11% discount rate. (If the net present value is negative, use either a negative sign preceding the number e.g. -45 or parentheses e.g. (45). Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 5,275.)


Net present value $Entry field with incorrect answer


Should the project be accepted?


The project should be Entry field with correct answer

accepted

.


LINK TO TEXT




Partially correct answer. Your answer is partially correct.  Try again.

   

Compute the net present value using the original estimates, but employing the 9% discount rate that Bob suggests is more appropriate. (If the net present value is negative, use either a negative sign preceding the number e.g. -45 or parentheses e.g. (45). Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 5,275.)


Net present value $Entry field with incorrect answer


Should the project be accepted?


The project should be Entry field with correct answer

accepted

.

Click if you would like to Show Work for this question:

Open Show Work
Answer #1

Part 1

YearPV factor @ 11%Remarks
1   0.90090= 1 / 1.11
2   0.81162= 0.9009 / 1.11
3   0.73119= 0.81162 / 1.11
4   0.65873= 0.73119 / 1.11
5   0.59345= 0.65873 / 1.11
6   0.53464= 0.59345 / 1.11
7   0.48166= 0.53464 / 1.11
8   0.43393= 0.48166 / 1.11
9   0.39092= 0.43393 / 1.11
10   0.35218= 0.39092 / 1.11
11   0.31728= 0.35218 / 1.11
12   0.28584= 0.31728 / 1.11
13   0.25751= 0.28584 / 1.11
14   0.23199= 0.25751 / 1.11
15   0.20900= 0.23199 / 1.11
Total   7.19087
Annual cash inflows         3,360,000

Less: Annual cash outflows           2,974,000

Net annual cash flows             386,000






AmountMultiply: PV factorPresent value
Cost of plant       (3,360,000)   1.00000          (3,360,000)
Net annual cash flows             386,000   7.19087             2,775,676
Salvage value         1,680,000   0.20900                351,120
Net present value

              (233,204)
This project should be rejected.

Rejected
[Negative NPV is rejected]


Part 2

Annual cash inflows (3360000+168000)         3,528,000

Less: Annual cash outflows (2974000-50000)         2,924,000

Net annual cash flows             604,000






AmountMultiply: PV factorPresent value
Cost of plant       (3,360,000)   1.00000          (3,360,000)
Net annual cash flows             604,000   7.19087             4,343,285
Salvage value         1,680,000   0.20900                351,120
Net present value

             1,334,405
This project should be Accepted

Accepted
[Positive NPV is accepted]


Part 3

YearPV factor @ 9%Remarks
1   0.91743= 1 / 1.09
2   0.84168= 0.91743 / 1.09
3   0.77218= 0.84168 / 1.09
4   0.70843= 0.77218 / 1.09
5   0.64993= 0.70843 / 1.09
6   0.59627= 0.64993 / 1.09
7   0.54703= 0.59627 / 1.09
8   0.50187= 0.54703 / 1.09
9   0.46043= 0.50187 / 1.09
10   0.42241= 0.46043 / 1.09
11   0.38753= 0.42241 / 1.09
12   0.35553= 0.38753 / 1.09
13   0.32618= 0.35553 / 1.09
14   0.29925= 0.32618 / 1.09
15   0.27454= 0.29925 / 1.09
Total   8.06069
Annual cash inflows         3,360,000

Less: Annual cash outflows           2,974,000

Net annual cash flows             386,000






AmountMultiply: PV factorPresent value
Cost of plant       (3,360,000)   1.00000          (3,360,000)
Net annual cash flows             386,000   8.06069             3,111,426
Salvage value         1,680,000   0.27454                461,227
Net present value

                212,654
This project should be Accepted

Accepted
[Positive NPV is accepted]



answered by: ANURANJAN SARSAM